A chemist has liquid fertilizer solutions that are 20% and 60% nitrogen. How many gallons of each of these solutions should the chemist mix to obtain 250 gallons of solution that is 30% nitrogen?
Answer by richwmiller(17219)

By substitution
We know the total amount is 250 gallons
We know the percentage of the total amount is 30%
We know the total amount has 75 gallons of 100%
We want to know the amount of solution at 20%
We want to know the amount of solution at 60%
x+y=250,
0.2*x+0.6*y=0.3*250
0.2*x+0.6*y=75
x=250-y
0.2*(250-y)+0.6*y=75
50-0.2y+0.6*y=75
0.4*y=25
We now know amount of solution at 60%
y=62.5 gallons at 60%
x=250-y
We now know amount of solution at 20%
x=187.5 gallons at 20%
check
0.2*187.5+0.6*62.5=0.3*250
37.5+37.5=75
75.0=75
ok